You have 4 KBytes of byte addressable memory. For this memory, you consider the following organizations: 1. 64-bit words, with a 9-bit address line. 2. 32-bit words, with an 11-bit address line. 3.16-bit words, with a 10-bit address line. In which organization(s) is it the case that not all words in the memory can be addressed?

a. 1 b. 2 c. 3 d. 2 & 3
Answer #1

Memory is 4KB = 12 bits

A) word size = 64 bit which is 8B

9 bits address line therefore size is 2^9*8B = 2^12 B

Therefore memory size is 12 bits

B) 32 bit word size = 4B word size

11 bit line. Therefore total memory size is 2^11 * 2^2 = 2^13B

Therefore memory size is 13 bit

C) 16 bit word is 2B word size

10 bit address line. Therefore total memory size is 2^10* 2= 2^11B

Therefore memory is 11 bits.

Correct option is (c). The third organization cannot address all the word. Since it's size is smaller than required memory size.

Please upvote... Thanks...

10 bit address line = 2^10*